# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 4.2 Using prompts
Figure 5 also depicts the task performance versus BERT layers. Our first observation is that for STS and clustering tasks, there is a strong preference for the first layers. As a result, the best combination of layers also involves the first ones: for semantic textual similarity tasks, it is the average of representations from 1, 2, and 12 layers; and for clustering, 0, 1, 2, and 12 layers. Even individual SICK-E, SICK-R, and STS-B classification tasks, which originate from semantic textual similarity ones, have strong first layers (see Figure 8 in Appendix A.2), although on average the best layer for classification tasks is the 10th. This shows that a lot of performance for tasks that work on similarities between texts (STS and clustering) depends on the first layers, which according to [234] have more generalized token representations. However, the last layer, which gets recreated token identities [234] is also important, as the best combinations (such as 1+2+12) also involve it. This first+last cooperation of layers can be distinctly observed as a U shape for clustering performance in layers (see Figure 5 and also Figure 4).
For most classification tasks (excluding those originating from STS), the U shape is inversed. This is very clearly seen for binary classification tasks (see Figure 8 in Appendix A.2), where the highest classification accuracy is concentrated between the 9th and the 11th layers. This aligns with the explanations of the previous work [100] which argues that the last layer is over-specialized for the training objective.
@@ 5 Conclusions
We empirically evaluated the effects of various aggregation and post-processing techniques of token representations in a trained transformer and other models to form good text-, paragraph-, or sentence-level embeddings. We did the empirical evaluation of the embeddings on three classes of downstream text-level tasks: Semantic Textual Similarity (STS), clustering, and classification.
@@
We present a strong and very simple baseline model of Random Embeddings (RE), where every token is assigned a random vector as its embedding. Combined with token aggregation and post-processing techniques, it almost matches the average STS performance of the BERT model, also with the techniques applied, with 66.4% versus 69.8% average Spearman correlation. It also shows very high performance for some tasks like stackoverflow classification (see Section 4.1.2 for more details). We encourage future work to use RE as a baseline, due to its mid-level performance, simple implementation, and ability to separate the contribution to the performance of the learned contexts from the aggregation and post-processing techniques.
We found that the aggregation and post-processing techniques tried typically increase the isotropy of the representations, and the isotropy for most models is positively correlated (up to 69% Pearson correlation) with the Spearman correlation of STS tasks (Section 4.1.3). The highest isotropy improvement is observed for our Random Embeddings model since its token representations have the maximal isotropy to start with. We did not find the token aggregation and post-processing techniques to improve the alignment and uniformity properties of representations.
@@
We presented a static vectors model Avg., which simply contains BERTs tokens averaged over multiple different contexts. Our experiments show that it outperforms a more complex BERT2Static [49], also a static word-level model yet specially trained on BERT contexts. With the best post-processing and token aggregation techniques, the advantage for unsupervised STS tasks is 69.5 versus 66.4, and for clustering it is 62.2 versus 57.2., with a negligible difference for classification tasks. Moreover, we show that combining Avg. with the parent BERT model can bring even further improvements. In particular, BERT+Avg. reached the highest average clustering accuracy of 64.8 out of all our considered models, as well as the classification accuracy of 80.5 (Section 4.3). We encourage future work to also use Avg. as a baseline, both due to its upper-level performance and simple implementation.
In our work, we also analyzed prompt and BERT+Avg. models layer-wise. We found that for the STS tasks taking the representations from the first layers is performing better, and for the clustering task, the performance profile forms a “U” shape with tops at the first and last layers. Therefore, for these two task groups we mostly use the average of first+last layers, harnessing both of the tops for the best performance. On the other hand, classification tasks have an inverted “U” shape with the top in the 10th layer (Figure 5). We did not find token aggregation and post-processing techniques to change such profile curvature.
In this research, we used a pre-trained BERT as a manageable representative of transformer models, but the findings should be transferable to other types of transformers, including large language models. We also specifically did not do task-specific fine-tuning of the model to keep it universal. This enables the same model to be used in multiple ways. For example, given a prompt, its text-level embedding can be extracted from the model using the techniques investigated here and this embedding can be used to search an external database for related information to add to the prompt to the same model; i.e., we can use the same model for both query encoding and generation in retrieval-augmented generation [235]. Alternatively, one of the simpler baseline models proposed here could be used as the query encoder.
@@ A.1 Token aggregation and post-processing
Table 6: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS12 task.
@@
Table 7: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS13 task.
@@
Table 8: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS14 task.
@@
Table 9: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS15 task.
@@
Table 10: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS16 task.
@@
Table 11: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS-B task.
@@
Table 12: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity SICK-R task.
@@
Table 13: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STR task.
@@
Table 14: Clustering accuracy dependence on token aggregation and post-processing techniques for the agnews dataset.
@@
Table 15: Clustering accuracy dependence on token aggregation and post-processing techniques for the biomedical dataset.
@@
Table 16: Clustering accuracy dependence on token aggregation and post-processing techniques for the googleTS dataset.
@@
Table 17: Clustering accuracy dependence on token aggregation and post-processing techniques for the searchsnippets dataset.
@@
Table 18: Clustering accuracy dependence on token aggregation and post-processing techniques for the stackoverflow dataset.
@@
Table 19: Clustering accuracy dependence on token aggregation and post-processing techniques for the tweet dataset.
@@
Figure 6: BERT+Avg. model individual task STS performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B+Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
Figure 7: BERT+Avg. model individual task clustering performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B+Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
Figure 8: BERT+Avg. model individual task supervised classification performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B+Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.

### v2

@@ 4.2 Using Prompts
Figure 5 also depicts the task performance versus BERT layers. Our first observation is that for STS and clustering tasks, there is a strong preference for the first layers. As a result, the best combination of layers also involves the first ones; for semantic textual similarity tasks, it is the average of representations from one, two, and twelve layers, and for clustering, zero, one, two, and twelve layers. Even individual SICK-E, SICK-R, and STS-B classification tasks, which originate from semantic textual similarity ones, have strong first layers (see Figure 8 in Appendix A.2), although on average, the best layer for classification tasks is the tenth. This shows that a lot of performance for tasks that work on similarities between texts (STS and clustering) depends on the first layers, which, according to [234], have more generalized token representations. However, the last layer, which receives recreated token identities [234] is also important, as the best combinations (such as 1 + 2 + 12) also involve it. This first + last cooperation of layers can be distinctly observed as a U shape for clustering performance in layers (see Figure 5 and also Figure 4).
For most classification tasks (excluding those originating from STS), the U shape is inversed. This is very clearly seen for binary classification tasks (see Figure 8 in Appendix A.2), where the highest classification accuracy is concentrated between the ninth and the eleventh layers. This aligns with the explanations of the previous work [100], which argues that the last layer is over-specialized for the training objective.
@@ 5 Conclusions
We empirically evaluated the effects of various aggregation and post-processing techniques of token representations in a trained transformer and other models to form good text-, paragraph-, or sentence-level embeddings. We carried out the empirical evaluation of the embeddings on three classes of downstream text-level tasks: Semantic Textual Similarity (STS), clustering, and classification.
@@
We present a strong and very simple baseline model of Random Embeddings (RE), where every token is assigned a random vector as its embedding. Combined with token aggregation and post-processing techniques, it also almost matches the average STS performance of the BERT model with the techniques applied, with 66.4% versus 69.8% average Spearman correlation. It also shows very high performance for some tasks, like stackoverflow classification, where BERT token contextualization may not work well on code samples in the texts (see Section 4.1.2 for more details). We encourage future work to use RE as a baseline, due to its mid-level performance, simple implementation, and ability to separate the contribution to the performance of the learned contexts from the aggregation and post-processing techniques.
We found that the aggregation and post-processing techniques tried typically increase the isotropy of the representations, and the isotropy for most models is positively correlated (up to 69% Pearson correlation) with the Spearman correlation of STS tasks (Section 4.1.3). The highest isotropy improvement is observed for our Random Embeddings model, since its token representations have the maximal isotropy to start with. We did not find the token aggregation and post-processing techniques to improve the alignment and uniformity properties of representations.
@@
We presented a static vector model Avg., which simply contains BERT tokens averaged over multiple different contexts. Our experiments show that it outperforms a more complex BERT2Static [49], also a static word-level model, yet specially trained on BERT contexts. With the best post-processing and token aggregation techniques, the advantage for unsupervised STS tasks is 69.5 versus 66.4, and for clustering, it is 62.2 versus 57.2., with a negligible difference for classification tasks. Moreover, we show that combining Avg. with the parent BERT model can bring even further improvements. In particular, BERT + Avg. reached the highest average clustering accuracy of 64.8 out of all our considered models, as well as the classification accuracy of 80.5 (Section 4.3). We encourage future work to also use Avg. as a baseline, both due to its upper-level performance and simple implementation.
In our work, we also analyzed prompt and BERT + Avg. models layer-wise. We found that for the STS tasks, taking the representations from the first layers performs better, and for the clustering task, the performance profile forms a “U” shape with tops at the first and last layers. Therefore, for these two task groups, we mostly use the average of first + last layers, harnessing both of the tops for the best performance. On the other hand, classification tasks have an inverted “U” shape with the top in the 10th layer (Figure 5). We did not find token aggregation and post-processing techniques to change such profile curvature.
In this research, we used a pre-trained BERT as a manageable representative of transformer models, but the findings should be transferable to other types of transformers, including large language models. We also specifically did not perform task-specific fine-tuning of the model to keep it universal. This enables the same model to be used in multiple ways. For example, given a prompt, its text-level embedding can be extracted from the model using the techniques investigated here, and this embedding can be used to search an external database for related information to add the prompt to the same model, i.e., we can use the same model for both query encoding and generation in retrieval-augmented generation [235]. Alternatively, one of the simpler baseline models proposed here could be used as the query encoder.
@@ A.1 Token Aggregation and Post-Processing
Table 6: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS12 task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 7: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS13 task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 8: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS14 task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 9: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS15 task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 10: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS16 task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 11: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STS-B task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 12: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity SICK-R task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 13: Spearman correlation dependence on token aggregation and post-processing techniques for semantic textual similarity STR task. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 14: Clustering accuracy dependence on token aggregation and post-processing techniques for the agnews dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 15: Clustering accuracy dependence on token aggregation and post-processing techniques for the biomedical dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 16: Clustering accuracy dependence on token aggregation and post-processing techniques for the googleTS dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 17: Clustering accuracy dependence on token aggregation and post-processing techniques for the searchsnippets dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 18: Clustering accuracy dependence on token aggregation and post-processing techniques for the stackoverflow dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Table 19: Clustering accuracy dependence on token aggregation and post-processing techniques for the tweet dataset. The best result for each model is bolded, while underlined result is the best across all the models.
@@
Figure 6: BERT + Avg. model individual task STS performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis, average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B + Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
Figure 7: BERT + Avg. model individual task clustering performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis, average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B + Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
Figure 8: BERT + Avg. model individual task supervised classification performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis, average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B + Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.